### v1

# LatentSync: Audio Conditioned Latent Diffusion Models for Lip Sync

Abstract
We present LatentSync, an end-to-end lip sync framework based on audio conditioned latent diffusion models without any intermediate motion representation, diverging from previous diffusion-based lip sync methods based on pixel space diffusion or two-stage generation. Our framework can leverage the powerful capabilities of Stable Diffusion to directly model complex audio-visual correlations. Additionally, we found that the diffusion-based lip sync methods exhibit inferior temporal consistency due to the inconsistency in the diffusion process across different frames. We propose Temporal REPresentation Alignment (TREPA) to enhance temporal consistency while preserving lip-sync accuracy. TREPA uses temporal representations extracted by large-scale self-supervised video models to align the generated frames with the ground truth frames. Furthermore, we observe the commonly encountered SyncNet convergence issue and conduct comprehensive empirical studies, identifying key factors affecting SyncNet convergence in terms of model architecture, training hyperparameters, and data preprocessing methods. We significantly improve the accuracy of SyncNet from 91% to 94% on the HDTF test set. Since we did not change the overall training framework of SyncNet, our experience can also be applied to other lip sync and audio-driven portrait animation methods that utilize SyncNet. Based on the above innovations, our method outperforms state-of-the-art lip sync methods across various metrics on the HDTF and VoxCeleb2 datasets. Code and models are publicly available at https://github.com/bytedance/LatentSync.
footnotetext: Corresponding authors
1 Introduction
The task of lip sync involves the generation of accurate lip movements in a talking video based on corresponding speech, while maintaining the head pose and personal identity. This technique has broad applications in numerous practical domains, such as dubbing, virtual avatars, and video conferencing.
Figure 1: Frameworks comparison between previous diffusion-based lip sync methods and our method.
In the field of lip sync, GAN-based methods [18, 36] remain the mainstream approachs. The main issue with GAN-based methods is that they struggle to scale up [24, 44] to large and diverse datasets due to the unstable training and mode collapse [47, 6]. Recent studies proposed diffusion-based methods [34, 2, 64, 57, 29] for lip sync, allowing the model to easily generalize across different individuals without the need for further fine-tuning on specific identities. However, these methods still have some limitations. Specifically, [34, 2] performs the diffusion process in the pixel space (Fig. 1 a), which restricts its ability to generate high-resolution videos due to the prohibitive hardware requirements. Methods like [64, 57] adopt a two-stage approach: the first stage generates lip motions from audio, and the second stage synthesizes the visual appearance conditioned on the motion (Fig. 1 b). The issue with this two-stage approach is that subtly different sounds may map to the same motion representation, which can result in the loss of nuanced expressions linked to the emotional tone of the speech.
To solve the above problems, we propose LatentSync, a novel end-to-end lipsync framework based on audio conditioned latent diffusion models [41], without the need of any intermediate 3D representation [3] or 2D landmarks [5, 30], as shown in Fig. 1 (c). LatentSync leverages the powerful generative capabilities of Stable Diffusion [41] to directly capture the intricate audio-visual correlations, facilitating the generation of dynamic and lifelike talking videos.
Furthermore, we found that the lip-sync accuracy of our model was quite low without SyncNet [9, 36] supervision. We explored two methods to add SyncNet supervision to latent diffusion models [41]: (a) adding SyncNet supervision after decoding to pixel space, and (b) training a SyncNet in the latent space. However, we discovered that SyncNet has convergence issues in both high-resolution pixel space and latent space. To solve this problem, we conducted comprehensive empirical studies in the aspects of model architecture, training hyperparameters, and data preprocessing methods, allowing the SyncNet to converge easily. We achieved the unprecedented 94% accuracy on HDTF [62].
Additionally, we found that in comparison to single-step lip sync methods [36, 34, 61, 18], diffusion-based lip sync approaches [34, 2], which employ multi-step generation processes, exhibit inferior temporal consistency. We hypothesize that this might be due to the inconsistency in the diffusion process across different frames. Inspired by [58, 15], we proposed a new method called TREPA, which uses a large-scale self-supervised video model VideoMAE-v2 [52] to extract a powerful representation rich in temporal information. The distance between the temporal representations of the generated consecutive frames and the ground truth consecutive frames is used as an extra loss. This method is very similar to LPIPS [59], with the key difference being that this loss leverages temporal representations. We found that this simple method effectively improves the temporal consistency while preserving the lip-sync accuracy of our model.
In summary, we make the following contributions: (1) We proposed an end-to-end lip sync framework based on audio conditioned latent diffusion model without any intermediate motion representation. (2) To solve the convergence issue of SyncNet, we conducted empirial studies in the aspects of model architecture, training hyperparameters, data preprocessing methods and identified the key factors affecting SyncNet convergence. (3) We proposed the TREPA to improve temporal consistency while preserving the lip-sync accuracy.
2 Related Work
2.1 Diffusion-based Lip Sync
[34] and [2] are both end-to-end lip sync methods based on pixel space audio conditioned diffusion models. [57] uses diffusion models in the first stage to complete audio-to-motion conversion and uses a VAE [26] in the second stage for motion-to-image generation. [64] uses transformers in the first stage to convert audio to motion and employs diffusion models in the second stage for motion-to-image generation. [29] uses diffusion autoencoders [37] to convert the masked images into semantic latent codes in the first stage, and uses diffusion models to generate an image conditioned on the semantic latent codes and audio in the second stage, which also belongs to pixel space diffusion models.
2.2 Non-diffusion-based Lip Sync
Wav2Lip [36] is the most classic lip sync method that introduced using a pretrained SyncNet [9] to supervise the training of the generator. [20] trains a VQ-VAE [12] to encode faces and head poses, and then trains the lip sync generator in the quantized space to generate high-resolution images. StyleSync [18] follows the overall framework of Wav2Lip, with its main innovation being the use of StyleGAN2 [25] as the generator backbone. [8] divides lip sync into three components: semantic-guided reenactment network, lip sync network, and identity-aware refinement and enhancement. DINet [63] deforms the feature maps to generate mouth shapes conditioned on the driving audio. MuseTalk [61] uses the architecture of Stable Diffusion [41] for inpainting, but it does not perform diffusion process and uses a discriminator for adversarial learning [17], making it more like a GAN-based framework.
3 Method
Figure 2: The overview of our LatentSync framework. We use the Whisper [39] to convert melspectrogram into audio embeddings, which are then integrated into the U-Net [42] via cross-attention layers. The reference and masked frames are channel-wise concatenated with noised latents as the input of U-Net. In the training process, we use one-step method to get estimated clean latents from predicted noises, which are then decoded to obtain the estimated clean frames. The TREPA, LPIPS [59] and SyncNet loss [36] are added in the pixel space.
3.1 LatentSync Framework
The overview of the LatentSync framework is shown in Fig. 2. Our backbone network generates the video frame by frame without temporal modeling between frames. The temporal consistency during the generation process relies entirely on the audio window. This frame-by-frame generation approach is beneficial for occlusion strategies in dubbing scenarios, as we can easily skip frames where the mouth is occluded and lip sync is not needed. Our framework is based on image-to-image inpainting [43], so it requires a masked image as input. Additionally, to integrate the visual features of the face from the original video, we need to input reference images. The reference images are randomly selected during training to avoid learning shortcuts [16], and they are taken from the current frames during inference. We concatenate these inputs along the channel dimension, so the total input consists of 13 channels (4 channels for noise latent, 1 channel for the mask, 4 channels for the masked image, and 4 channels for the reference image). At the beginning of training, we initialized our model with the parameters of SD 1.5 [41], except for the first conv_in layer with 13 channels and cross-attention layers of dimension 384, which are randomly initialized. Theoretically, our framework is compatible with more advanced latent diffusion models such as SDXL [35], SD3 [13], and FLUX [28], but we only conducted experiments on SD 1.5 due to hardware limitations.
Audio layers. We used the pretrained audio feature extractor Whisper [39] to extract audio embeddings. Lip motion may be influenced by the audio from the surrounding frames, and a larger range of audio input also provides more temporal information for the model. Therefore, for each generated frame, we bundled the audio from several surrounding frames as input. We define the input audio feature $A^{(f)}$ for the $f$th frame as: $A^{(f)}=\left\{a^{(f-m)},\dots,a^{(f)},\dots,a^{(f+m)}\right\}$, where $m$ is the number of surrounding audio features from one side. To integrate the audio embeddings into the U-Net [42], we used the native cross-attention layer.
SyncNet supervision. Our model predicts in the noise space, while SyncNet [9, 36] requires an input in the image space. To address this issue, we use the predicted noise $\epsilon_{\theta}(z_{t})$ to obtain the estimated $\hat{z}_{0}$ in one step, which can be formulated as:
$\hat{z}_{0}=\left(z_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(z_{t})\right)/\sqrt{\bar{\alpha}_{t}}$ (1)
Another problem is that our model makes predictions in the latent space. We explored two methods to add SyncNet supervision to latent diffusion models: (a) Decoded pixel space supervision, which trains SyncNet in the same way as [36]. (b) Latent space supervision, which requires training a SyncNet in the latent space. The visual encoder input of this SyncNet is the latent vectors obtained by the VAE [12, 26] encoding. The illustration is shown in Fig. 3. We found that training SyncNet in the latent space converges worse compared to training it in the pixel space. We speculate that this is due to information loss in the lip area during the VAE encoding process. The poorer convergence of the latent space SyncNet also results in the worse audio-lip sync accuracy for the supervised diffusion models, according to the experimental results in Sec. 4.3. Therefore, we use decoded pixel space supervision in the LatentSync framework.
Figure 3: Two methods to add SyncNet supervision to latent diffusion models.
Two-stage training. In the first stage, we do not decode to the pixel space and do not add SyncNet loss. The purpose of the first stage is to allow the U-Net to learn visual features using a large batch size. Therefore, the training objective of the first stage has only a simple loss [23]:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x,A,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(A))\right\|_{2}^{2}\right]$ (2)
where $A$ is the input audio, $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(A))$ is the predicted noise, and $\tau_{\theta}$ is the audio feature extractor.
In the second stage of training, we used the decoded pixel space supervision method to add SyncNet loss. Compared to [36], we increased the input frame length of SyncNet to 16. For detailed changes to the SyncNet structure, please refer to Sec. 3.3. By using the 16 decoded video frames $\mathcal{D}(\hat{z}_{0})_{f:f+16}$ and the corresponding audio sequence $a_{f:f+16}$, the SyncNet loss can be formulated as:
$\mathcal{L}_{\text{sync}}=\mathbb{E}_{x,a,\epsilon,t}\left[\text{SyncNet}(\mathcal{D}(\hat{z}_{0})_{f:f+16},a_{f:f+16})\right]$ (3)
where $\mathcal{D}$ represents the VAE decoder, since the lipsync task requires generating detailed areas, such as lips, teeth, and facial hair, we used the LPIPS [59] to improve the visual quality of the images generated by the U-Net.
$\mathcal{L}_{\text{lpips}}=\mathbb{E}_{x,\epsilon,t}\left[\left\|\mathcal{V}_{l}(\mathcal{D}(\hat{z}_{0})_{f})-\mathcal{V}_{l}(x_{f})\right\|_{2}^{2}\right]$ (4)
where $\mathcal{V}_{l}(\cdot)$ denotes the features extracted from the $l^{\text{th}}$ layer of a pretrained VGG network [45]. In addition, to improve temporal consistency, we also employed the proposed TREPA, see Eq. 7. For more details of the TREPA, please refer to Sec. 3.2.
The total loss function for the second stage of training is:
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{simple}}+\lambda_{2}\mathcal{L}_{\text{sync}}+\lambda_{3}\mathcal{L}_{\text{lpips}}+\lambda_{4}\mathcal{L}_{\text{trepa}}$ (5)
We identified the optimal coefficients through empirical analysis as follows: $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})=(1,0.05,0.1,10)$
Mixed noise model. To ensure that the model correctly learns temporal information, the input noises also need to have temporal consistency; otherwise the model will fail to learn properly. The simplest method is applying the same noise to all frames [34]. We further found that diffusion models trained with the mixed noise model [14] exhibit better temporal consistency. The method generates two distinct noise vectors, $\epsilon_{\text{shared}}$ and $\epsilon_{\text{ind}}$. $\epsilon_{\text{shared}}$ represents a common noise vector applied across all 16 video frames, whereas $\epsilon_{\text{ind}}$ corresponds to the frame-specific noise. The final noise is obtained through a linear combination of these two vectors.
$\epsilon^{f}=\epsilon_{\text{shared}}+\epsilon_{\text{ind}}^{f}$ (6)
We used the mixed noise model in both two training stages, and we applied the same initial latent noise to all frames in the video during inference stage.
Affine transformation. During the data preprocessing stage, affine transformation was employed to achieve face frontalization. This approach [18] helps the model to effectively learn facial features particularly in challenging scenarios such as side-profile views. We applied a mask that covers the entire face to prevent information leakage, ensuring that the model does not infer lip movements based on facial movements. The position and shape of the mask are fixed. We do not use detected landmarks [30, 5] to draw the mask as they will leak information about facial movements. The affine transformation and fixed mask are illustrated in Fig. 4.
Figure 4: The illustration of affine transformation and fixed mask.
3.2 Temporal Representation Alignment
TREPA aligns the temporal representations of the generated image sequences with those of ground truth image sequences. The insight behind this method is that merely employing distance loss between individual images only improves the content quality of single generated images but does not enhance the temporal consistency of the generated image sequence. In contrast, temporal representations capture temporal correlation within image sequences, enabling the model to focus on improving the overall temporal consistency. We employed a large-scale self-supervised video model VideoMAE-v2 [52] to extract temporal representations. Due to its unsupervised training on large-scale unlabeled datasets, the model’s temporal representations exhibit strong generalization capabilities, robustness, and high information density.
In mathematical form, let $\mathcal{T}$ be the self-supervised video model encoder. The encoder’s output is the embedding before the head projection. TREPA can be represented as:
$\mathcal{L}_{\text{trepa}}=\mathbb{E}_{x,\epsilon,t}\left[\left\|\mathcal{T}(\mathcal{D}(\hat{z}_{0})_{f:f+16})-\mathcal{T}(x_{f:f+16})\right\|_{2}^{2}\right]$ (7)
where the straightforward Mean Squared Error (MSE) is employed to measure the distance between temporal representations. We also fix the representation by $\ell_{2}$ normalization before calculating the MSE.
Observations. The temporal layer [19, 4] adds temporal self-attention layers to the U-Net to enhance temporal consistency. We initially attempted to apply the temporal layer to LatentSync, but found that it significantly compromised lip-sync accuracy. In contrast, our TREPA not only avoids harming lip-sync accuracy but even improves it to some extent. We hypothesize this is because the temporal layer adds additional parameters to the U-Net, causing some of the gradients in backpropagation to be distributed to the temporal layer parameters, weakening the learning of the audio cross-attention layer parameters. On the other hand, TREPA does not add extra parameters; to improve temporal consistency, the model must rely more effectively on the information in the audio window (since we know that lip-sync models use the audio window to capture temporal information). As a result, the audio cross-attention layers are further trained and strengthened. Refer to the quantitative results in Sec. 4.3.
Another interesting finding is that TREPA can also be applied to improve the temporal consistency and lip-sync accuracy of Wav2Lip [36], according to the experimental results in Sec. 4.3. Theoretically, any video generation model that generates multiple consecutive frames at once can utilize TREPA to enhance the temporal consistency.
3.3 Empirical Studies on SyncNet Convergence
Many researchers have reported the convergence problem of SyncNet on GitHub issues [32, 56, 40, 33], a typical characteristic of this issue is that the training loss gets stuck at 0.69 and fails to decrease further. However, no academic work has provided a detailed analysis of this problem to date. In this section, we analyze the SyncNet convergence problem and identify several critical factors affecting the convergence of SyncNet through various ablation studies. Importantly, we preserved SyncNet’s original training framework and employed the same contrastive loss function as the original paper [9]. Therefore, our experience can be applied to many lip sync [36, 18, 34] and audio-driven portrait animation methods [31] that utilize SyncNet.
Why is the SyncNet training loss stuck at 0.69? According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. When $q(x=1)\approx q(x=0)\approx 0.5$ and batch size $N$ is sufficiently large, the training loss of SyncNet is (proof in Appendix B):
$\mathcal{L}_{\text{syncnet}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\sum_{x_{i}\in\{0,1\}}p(x_{i})\,\text{log}\,{q(x_{i})}\approx 0.693$ (8)
This means that SyncNet has not learned any discriminative capability; it is just randomly guessing whether the samples are positive or negative. This may be due to various reasons, including the model’s insufficient capacity to fit the data, the large audio-visual offset in the data, and flaws in the training strategy. In the following paragraphs, we will identify the key factors affecting the convergence of SyncNet through comprehensive ablation studies.
Batch size. As shown in Fig. 5, a larger batch size (e.g., 1024) not only enables the model to converge faster and more stably but also results in a lower validation loss at the end of training. In contrast, smaller batch sizes (e.g., 128) may fail to converge, with the loss remaining stuck at 0.69. Even with a slightly larger batch size (e.g., 256), while convergence may be achieved, the training loss exhibits significant oscillations during its descent.
Figure 5: SyncNet training curves of different batch sizes. VoxCeleb2 results, the more transparent curves represent the training set loss, while the darker curves represent the validation set loss; the same applies to the following figures. (VoxCeleb2, Dim 2048, SD residual blocks, 16 frames.)
Architecture. We attempted to replace the SyncNet’s visual and audio encoders with the U-Net encoder from SD 1.5 [41], retaining the structure of residual blocks [21] and self-attention blocks [51] in the U-Net encoder blocks. We only adjusted the downsampling factors based on the size of the input visual images and mel-spectrograms, and we removed the cross-attention blocks, as SyncNet does not require additional conditions. As shown in Fig. 6, SyncNet using the SD U-Net encoder architecture maintained both training loss and validation loss lower than those of Wav2Lip’s SyncNet [36] throughout the training process, though the difference was not substantial.
Figure 6: SyncNet training curves of different architectures. For comparison, we also modified the architecture of Wav2Lip’s SyncNet to accept 256 $\times$ 256 visual input according to [38]. (VoxCeleb2, Dim 2048, Batch size 512, 5 frames.)
Self attention. We attempted to remove the self-attention [51] blocks in the U-Net encoder, keeping only the residual blocks [21]. As shown in Fig. 7, removing the self-attention blocks results in faster convergence during the initial training phase, while achieving lower validation loss towards the end of the training. Additionally, this approach helps to reduce the model size and accelerates both inference and training speeds. We hypothesize that the reason may be that attention layers do not possess the inherent inductive biases of CNNs, such as translation equivariance and locality. Therefore, they do not generalize well without sufficient data. Similar findings and explanations can be found in [11].
Figure 7: SyncNet training curves with and without self-attention layers. (VoxCeleb2, Dim 2048, Batch size 512, 16 frames.)
Number of frames. The number of input frames determines the range of visual and audio information that SyncNet can perceive. As shown in Fig. 8, selecting a larger number of frames (e.g., 16) can help the model converge. However, an excessively large number of frames (e.g., 25) will cause the model to get stuck around 0.69 in the early stage of training, and the validation loss at the end of training does not show a significant advantage compared to using 16 frames.
Figure 8: SyncNet training curves of different numbers of input frames. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048.)
Data preprocessing. In-the-wild videos naturally contain audio-visual offsets, so it is necessary to adjust this offset to zero before inputting them into the SyncNet network. We used the official open-source version of SyncNet [9] to adjust the offset and remove videos with ${\text{Sync}}_{\text{conf}}$ below 3. Specifically, we tested adjusting the offset before and after applying affine transformation. As shown in Fig. 9, without offset adjustment, the model’s convergence is significantly impaired. Performing affine transformation before offset adjustment yields better results. We hypothesize that this may be due to affine transformation reducing data with side profiles or unusual angles, allowing the official SyncNet [9] to predict the offset more accurately.
Figure 9: SyncNet training curves of different data preprocessing methods. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048, 16 frames.)
Discussions. Based on the ablation studies above, we found that batch size, number of input frames, and data preprocessing method significantly impact SyncNet convergence. The model architecture can also affect convergence, though less noticeably. We identified the optimal settings for a SyncNet with the 256 $\times$ 256 input size: batch size of 1024, 16 frames, SD residual blocks for both visual and audio encoders, embedding dimension of 2048, and adjusting the audio-visual offset after affine transformation. We train a SyncNet on VoxCeleb2 [10] and test it on HDTF [62], which is an out-of-distribution experimental setup. The validation loss on VoxCeleb2 reaches around 0.18 and the accuracy on HDTF achieves 94%, which significantly surpasses the previous SOTA result 91% [36, 20]. For more ablation studies, please refer to Appendix C.
4 Experiments
Method HDTF VoxCeleb2
FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
Wav2Lip [36] 12.5 0.70 8.2 304.35 10.8 0.71 7.0 257.85
VideoReTalking [8] 9.5 0.75 7.5 270.56 7.5 0.77 6.4 215.67
DINet [63] 8.02 0.79 7.7 208.93 6.3 0.80 6.7 110.75
MuseTalk [61] 9.35 0.74 6.8 246.75 7.1 0.80 5.9 203.43
LatentSync (Ours) 7.03 0.79 8.9 192.74 5.6 0.81 7.3 124.38
Table 1: Quantitative comparisons on HDTF and VoxCeleb2.
4.1 Experimental Settings
Datasets. We used a mixture of VoxCeleb2 [10] and HDTF [62] datasets as our training set. VoxCeleb2 is a large-scale audio-visual dataset containing over 1 million utterances from over 6,000 speakers. It includes speakers from a wide range of ethnicities, accents, and backgrounds. HDTF contains 362 different high-definition (HD) videos, with resolutions typically around 720p to 1080p.
We used HyperIQA [48] to filter out videos with low visual quality, specifically blurry or pixelated videos. During evaluation, we randomly selected 30 videos from the test set of HDTF or VoxCeleb2.
Implementation details. When evaluating our model LatentSync, we first converted the videos to 25 FPS, then applied the affine transformation based on facial landmarks detected by face-alignment [5] to obtain 256 $\times$ 256 face videos. The audio was resampled to 16kHz. We used 20 steps of DDIM [46] sampling for inference.
Evaluation metrics. We evaluate our method in three aspects: (1) Visual quality. We use SSIM [53] in the reconstruction setting and FID [22] in the cross generation setting to assess visual quality. Following [34], the reconstruction setting refers to using the same audio as the input video, while the cross generation setting refers to using audio different from the input video. (2) Lip-sync accuracy. We use the confidence score of SyncNet (${\text{Sync}}_{\text{conf}}$) [9], which we found to align closely with visual assessments. (3) Temporal consistency. We adopt the widely used FVD metric [50].
4.2 Comparisons
Comparison methods. We selected several SOTA methods that provide open-source inference code and checkpoints for comparison. Wav2Lip [36] is the classic lip-sync method, introducing the idea of using a pretrained SyncNet for supervision instead of a lip-sync discriminator [27]. VideoReTalking [8] divides the lip-sync process into three steps to improve the results. DINet [63] proposed performing deformations on feature maps to achieve high-quality lip sync. MuseTalk [61] utilizes the architecture of Stable Diffusion for inpainting, but it is not based on the diffusion model framework and appears more like a GAN-based approach.
Quantitative comparisons. As shown in Tab. 1, the lip-sync accuracy of our method significantly surpasses that of other methods. We believe this may be attributed to our 94% accuracy SyncNet, as well as the audio cross-attention layers, which may better capture the relationship between audio and lip movements. In terms of visual quality, our method outperforms others, likely due to the capabilities of the Stable Diffusion model. Owing to the incorporation of TREPA, our FVD score is also superior to other methods.
Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
LatentSync w/o SyncNet 1.6 260.37
LatentSync + latent space SyncNet 7.9 240.45
LatentSync + pixel space SyncNet 8.9 192.74
Table 2: The ablation studies of different SyncNets. HDTF results.
Qualitative comparisons. We run two cases in the cross generation setting. According to Fig. 10, Wav2Lip has excellent lip-sync accuracy, but the generated videos are very blurry. VideoReTalking shows strange artifacts. DINet’s generated box is obvious, and its mouth movements are stiff. MuseTalk does not preserve facial features well, the man’s beard becomes sparse. In contrast, our method excels in both clarity and identity preservation, even the mole on the woman’s face was preserved. Furthermore, it hardly shows generated box due to the smooth shape of fixed mask.
Figure 10: Qualitative comparisons with SOTA lip-sync methods. We run two cases in the cross generation setting [34]. The first row demonstrates the original input video, and the second row is the video from which we extracted the audio as input, the video can be regarded as the target lip movements. Rows 3 $\sim$ 7 display the lip-synced videos. (All the photorealistic portrait images in this paper are from licensed models.)
Figure 11: SyncNet training curves of different visual input spaces. Here we use the open-sourced pretrained VAE from Stability AI [1] for encoding. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 1024, 16 frames.)
4.3 Ablation studies
The supervision effect of SyncNet. As shown in Tab. 2, when the SyncNet supervision is not added, the trained diffusion models have no lipsync ability at all. In fact, we found that other end-to-end lipsync methods [36, 34] exhibit similar phenomena. As for supervision in two different spaces, diffusion models under pixel space SyncNet supervision perform better in terms of lip-sync accuracy and temporal consistency. This is likely due to the poor convergence of latent space SyncNet supervision (as shown in Fig. 11), which is quite reasonable since the input to the latent space SyncNet is the compressed latents obtained from VAE encoding, and some lip information may already be lost.
In addition, we found that improvements in lip-sync accuracy were accompanied by an increase in temporal consistency. This is because the temporal consistency of our model is provided by the audio window; as the model becomes better at utilizing audio information, temporal consistency also improves accordingly.
Different methods for improving temporal consistency. According to Tab. 3, the temporal layer significantly improves temporal consistency but also noticeably harms lip-sync accuracy. Our TREPA achieves balanced improvements in both lip-sync accuracy and temporal consistency without evident drawbacks. Considering the importance of lip-sync accuracy, TREPA is obviously the better approach.
Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
LatentSync 8.2 234.35
LatentSync + Temporal layer [19] 4.7 174.56
LatentSync + TREPA 8.9 192.74
Table 3: The ablation studies of different methods for improving temporal consistency. HDTF results.
In addition, TREPA can also improve Wav2Lip according to Tab. 4, which suggests that our method has general applicability in lip-sync tasks.
Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
Wav2Lip 8.2 304.35
Wav2Lip + TREPA 8.6 264.56
Table 4: The ablation studies of Wav2Lip with TREPA. HDTF results.
5 Conclusion
We introduced a new lip sync framework named LatentSync that addresses the problems of traditional diffusion-based lip sync methods: (1) The low sampling speed and inability to generate high-resolution videos of pixel space diffusion. (2) The information loss in two-stage generation methods. In addition, we proposed TREPA to improve temporal consistency while preserving the lip-sync accuracy, which mitigates the problem of video flickering in diffusion-based lip sync methods. Furthermore, we analyzed the notorious SyncNet convergence issue and conducted comprehensive ablation studies to identify several key influencing factors.
Appendix
Appendix A The Differences between Lip Sync and Audio-driven Portrait Animation
Many people may confuse lip sync with audio-driven portrait animation. These two tasks have some similarities but are actually completely different tasks. Their main difference lies in the input:
  • The inputs of Lip Sync: Video + Audio
  • The inputs of Audio-driven Portrait Animation: Image + Audio
This leads to differences in their overall frameworks. Lip sync is more like a video-to-video editing framework, which needs to keep areas other than the mouth the same as in the input video. Audio-driven portrait animation is more like an image-to-video animation framework, which can change the head’s movement and even facial expressions. The differences in the overall frameworks result in substantial differences in their training and inference methods. We list some methods for the two different tasks to help readers better distinguish between them:
  • Lip Sync: Wav2Lip[36], Diff2Lip [34], StyleSync [18], MuseTalk [61], MyTalk [57]
  • Audio-driven Portrait Animation: EMO [49], Hallo [54], EchoMimic [7], VASA-1 [55], DreamTalk [31], SadTalker [60]
Although there are already some audio-driven portrait animation methods based on Stable Diffusion [41], such as EMO [49] and Hallo [54], it is hard to adapt them for the lip sync task. For example, the temporal layer is very effective in these methods, but we found that it is not suitable for the lip sync task. To the best of our knowledge, the methods closest to ours are Diff2Lip [34] and MyTalk [57], which are pixel space diffusion and two-stage generation methods.
Appendix B SyncNet Training Loss Proof
According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage. The output of SyncNet is a probability distribution of whether the sample is positive or negative. We define $p(x=1)$ as the probability that the sample is positive, and $p(x=0)$ as the probability that the sample is negative. Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution. We plot some scatter charts to observe the changes in the prediction probabilities of a non-converging SyncNet during the training process.
Figure 12: We train a SyncNet on VoxCeleb2 [10] with batch size of 256. We modify its architecture and data to make it non-converging. Every 100 steps, we plot a scatter plot to observe the probability distribution of SyncNet’s predicted $q(x_{i}=1)$, including 256 data points. The x-axis represents the probability $q(x_{i}=1)$, and we add some random jitter along the y-axis for better visualization.
As shown in Fig. 12, in the early stages of training, the probabilities predicted by SyncNet for $q(x_{i}=1)$ are dispersed, but later in the training, the probabilities for $q(x_{i}=1)$ are almost all distributed around 0.5 (the scatter charts for $q(x_{i}=0)$ are similar). Now we have:
$\forall i\in{1,2,\ldots,N}:~{}~{}q(x_{i}=1)\approx q(x_{i}=0)\approx 0.5$ (9)
We speculate that this may be because, the non-converging SyncNet is consistently unable to correctly distinguish whether a sample is positive or negative during the training process. However, to reduce the training loss, it tends to output a probability of 0.5 for all input embeddings to minimize the training loss as much as possible. This is why the training loss decreases somewhat initially and then gets stuck (from 0.8 to 0.69).
We assmue that there are $A$ positive samples and $B$ negative samples in a batch, and then we can derive SyncNet’s training loss based on the following steps:
$\mathcal{L}_{\text{syncnet}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\sum_{x_{i}\in\{0,1\}}p(x_{i})\,\text{log}\,{q(x_{i})}$ (10)
$=-\frac{1}{N}\sum_{i=1}^{N}\left[p(x_{i}=1)\,\text{log}\,{q(x_{i}=1)}+p(x_{i}=0)\,\text{log}\,{q(x_{i}=0)}\right]$ (11)
$=-\frac{1}{N}\sum_{i=1}^{N}p(x_{i}=1)\,\text{log}\,{q(x_{i}=1)}-\frac{1}{N}\sum_{i=1}^{N}p(x_{i}=0)\,\text{log}\,{q(x_{i}=0)}$ (12)
$=-\frac{1}{N}\left[\sum_{a=1}^{A}1\times\,\text{log}\,{q(x_{a}=1)}+\sum_{b=1}^{B}0\times\,\text{log}\,{q(x_{b}=1)}\right]-\frac{1}{N}\left[\sum_{a=1}^{A}0\times\,\text{log}\,{q(x_{a}=1)}+\sum_{b=1}^{B}1\times\,\text{log}\,{q(x_{b}=1)}\right]$ (13)
$=-\frac{1}{N}\left[\sum_{a=1}^{A}\,\text{log}\,{q(x_{a}=1)}+\sum_{b=1}^{B}\,\text{log}\,{q(x_{b}=0)}\right]$ (14)
[…]
$=0.693$ (16)
Appendix C Additional SyncNet Ablation Studies
C.1 Embedding dimension
The dimensions of the audio and visual embeddings in SyncNet should be carefully calibrated to a moderate size. As illustrated in Fig. 13, embeddings with smaller dimensions (e.g., 512) result in representations that fail to capture sufficient semantic information, while larger dimensions (e.g., 4096 or 6144) lead to sparse representations, thereby impeding model convergence. Through empirical analysis, we identified that an optimal embedding dimension of 2048 is suitable for images with an input resolution of 256 $\times$ 256 pixels. For images of varying resolutions, the embedding dimensions should be appropriately scaled to ensure optimal performance.
Figure 13: SyncNet training curves of different embedding dimensions. (VoxCeleb2, Batch size 512, SD residual blocks, 16 frames.)

### v2

@@
-LatentSync: Audio Conditioned Latent Diffusion Models for Lip Sync
+LatentSync: Taming Audio-Conditioned Latent Diffusion Models for Lip Sync with SyncNet Supervision
 Abstract
-We present LatentSync, an end-to-end lip sync framework based on audio conditioned latent diffusion models without any intermediate motion representation, diverging from previous diffusion-based lip sync methods based on pixel space diffusion or two-stage generation.
-Our framework can leverage the powerful capabilities of Stable Diffusion to directly model complex audio-visual correlations.
-Additionally, we found that the diffusion-based lip sync methods exhibit inferior temporal consistency due to the inconsistency in the diffusion process across different frames.
-We propose Temporal REPresentation Alignment (TREPA) to enhance temporal consistency while preserving lip-sync accuracy.
-TREPA uses temporal representations extracted by large-scale self-supervised video models to align the generated frames with the ground truth frames.
-Furthermore, we observe the commonly encountered SyncNet convergence issue and conduct comprehensive empirical studies, identifying key factors affecting SyncNet convergence in terms of model architecture, training hyperparameters, and data preprocessing methods.
-We significantly improve the accuracy of SyncNet from 91% to 94% on the HDTF test set.
-Since we did not change the overall training framework of SyncNet, our experience can also be applied to other lip sync and audio-driven portrait animation methods that utilize SyncNet.
-Based on the above innovations, our method outperforms state-of-the-art lip sync methods across various metrics on the HDTF and VoxCeleb2 datasets.
+End-to-end audio-conditioned latent diffusion models (LDMs) have been widely adopted for audio-driven portrait animation, demonstrating their effectiveness in generating lifelike and high-resolution talking videos.
+However, direct application of audio-conditioned LDMs to lip-synchronization (lip-sync) tasks results in suboptimal lip-sync accuracy.
+Through an in-depth analysis, we identified the underlying cause as the “shortcut learning problem”, wherein the model predominantly learns visual-visual shortcuts while neglecting the critical audio-visual correlations.
+To address this issue, we explored different approaches for integrating SyncNet supervision into audio-conditioned LDMs to explicitly enforce the learning of audio-visual correlations.
+Since the performance of SyncNet directly influences the lip-sync accuracy of the supervised model, the training of a well-converged SyncNet becomes crucial.
+We conducted the first comprehensive empirical studies to identify key factors affecting SyncNet convergence.
+Based on our analysis, we introduce StableSyncNet, with an architecture designed for stable convergence.
+Our StableSyncNet achieved a significant improvement in accuracy, increasing from 91% to 94% on the HDTF test set.
+Additionally, we introduce a novel Temporal Representation Alignment (TREPA) mechanism to enhance temporal consistency in the generated videos.
+Experimental results show that our method surpasses state-of-the-art lip-sync approaches across various evaluation metrics on the HDTF and VoxCeleb2 datasets.
 Code and models are publicly available at https://github.com/bytedance/LatentSync.
-footnotetext: Corresponding authors
 1 Introduction
-The task of lip sync involves the generation of accurate lip movements in a talking video based on corresponding speech, while maintaining the head pose and personal identity.
-This technique has broad applications in numerous practical domains, such as dubbing, virtual avatars, and video conferencing.
-Figure 1: Frameworks comparison between previous diffusion-based lip sync methods and our method.
+The lip sync [29, 28, 14, 52] is a video editing task, which regenerates the lip movements of a talking person according to the given audio, while maintaining the head pose and personal identity.
+This technique has broad applications in numerous practical domains, such as visual dubbing, virtual avatars, and video conferencing.
+Figure 1: Frameworks comparison between previous diffusion-based lip-sync methods and our method.
 In the field of lip sync, GAN-based methods [14, 29] remain the mainstream approachs.
-The main issue with GAN-based methods is that they struggle to scale up [24, 44] to large and diverse datasets due to the unstable training and mode collapse [47, 6].
+The main issue with these methods is that they struggle to scale up [21, 36] to large and diverse datasets due to the unstable training [2, 35] and mode collapse [39, 6].
 Recent studies proposed diffusion-based methods [28, 3, 53, 48, 25] for lip sync, allowing the model to easily generalize across different individuals without the need for further fine-tuning on specific identities.
 However, these methods still have some limitations.
-Specifically, [34, 2] performs the diffusion process in the pixel space (Fig.
+Specifically, [28, 3] perform the diffusion process in the pixel space (Fig.
 1 a), which restricts its ability to generate high-resolution videos due to the prohibitive hardware requirements.
-Methods like [64, 57] adopt a two-stage approach: the first stage generates lip motions from audio, and the second stage synthesizes the visual appearance conditioned on the motion (Fig.
+Other methods [53, 48] adopt a two-stage approach: the first stage generates lip motions from audio, and the second stage synthesizes the visual appearance conditioned on the motion (Fig.
 1 b).
-The issue with this two-stage approach is that subtly different sounds may map to the same motion representation, which can result in the loss of nuanced expressions linked to the emotional tone of the speech.
-To solve the above problems, we propose LatentSync, a novel end-to-end lipsync framework based on audio conditioned latent diffusion models [41], without the need of any intermediate 3D representation [3] or 2D landmarks [5, 30], as shown in Fig.
+The issue with this two-stage approach is that subtly different sounds may map to the same motion representation, leading to the loss of nuanced expressions linked to the emotional tone of the speech.
+To address the above limitations, we propose LatentSync, an end-to-end lip sync framework based on audio-conditioned LDMs [33] to generate lifelike and high-resolution talking videos, as shown in Fig.
 1 (c).
-LatentSync leverages the powerful generative capabilities of Stable Diffusion [41] to directly capture the intricate audio-visual correlations, facilitating the generation of dynamic and lifelike talking videos.
-Furthermore, we found that the lip-sync accuracy of our model was quite low without SyncNet [9, 36] supervision.
-We explored two methods to add SyncNet supervision to latent diffusion models [41]: (a) adding SyncNet supervision after decoding to pixel space, and (b) training a SyncNet in the latent space.
-However, we discovered that SyncNet has convergence issues in both high-resolution pixel space and latent space.
-To solve this problem, we conducted comprehensive empirical studies in the aspects of model architecture, training hyperparameters, and data preprocessing methods, allowing the SyncNet to converge easily.
-We achieved the unprecedented 94% accuracy on HDTF [62].
-Additionally, we found that in comparison to single-step lip sync methods [36, 34, 61, 18], diffusion-based lip sync approaches [34, 2], which employ multi-step generation processes, exhibit inferior temporal consistency.
-We hypothesize that this might be due to the inconsistency in the diffusion process across different frames.
-Inspired by [58, 15], we proposed a new method called TREPA, which uses a large-scale self-supervised video model VideoMAE-v2 [52] to extract a powerful representation rich in temporal information.
-The distance between the temporal representations of the generated consecutive frames and the ground truth consecutive frames is used as an extra loss.
-This method is very similar to LPIPS [59], with the key difference being that this loss leverages temporal representations.
-We found that this simple method effectively improves the temporal consistency while preserving the lip-sync accuracy of our model.
-In summary, we make the following contributions: (1) We proposed an end-to-end lip sync framework based on audio conditioned latent diffusion model without any intermediate motion representation.
-(2) To solve the convergence issue of SyncNet, we conducted empirial studies in the aspects of model architecture, training hyperparameters, data preprocessing methods and identified the key factors affecting SyncNet convergence.
-(3) We proposed the TREPA to improve temporal consistency while preserving the lip-sync accuracy.
-2 Related Work
-2.1 Diffusion-based Lip Sync
-[34] and [2] are both end-to-end lip sync methods based on pixel space audio conditioned diffusion models.
-[57] uses diffusion models in the first stage to complete audio-to-motion conversion and uses a VAE [26] in the second stage for motion-to-image generation.
-[64] uses transformers in the first stage to convert audio to motion and employs diffusion models in the second stage for motion-to-image generation.
-[29] uses diffusion autoencoders [37] to convert the masked images into semantic latent codes in the first stage, and uses diffusion models to generate an image conditioned on the semantic latent codes and audio in the second stage, which also belongs to pixel space diffusion models.
-2.2 Non-diffusion-based Lip Sync
-Wav2Lip [36] is the most classic lip sync method that introduced using a pretrained SyncNet [9] to supervise the training of the generator.
-[20] trains a VQ-VAE [12] to encode faces and head poses, and then trains the lip sync generator in the quantized space to generate high-resolution images.
-StyleSync [18] follows the overall framework of Wav2Lip, with its main innovation being the use of StyleGAN2 [25] as the generator backbone.
-[8] divides lip sync into three components: semantic-guided reenactment network, lip sync network, and identity-aware refinement and enhancement.
-DINet [63] deforms the feature maps to generate mouth shapes conditioned on the driving audio.
-MuseTalk [61] uses the architecture of Stable Diffusion [41] for inpainting, but it does not perform diffusion process and uses a discriminator for adversarial learning [17], making it more like a GAN-based framework.
-3 Method
+We initially tried directly applying methods from the field of audio-driven portrait animation, such as EMO [41] and Hallo [46].
+However, the results showed poor lip-sync accuracy.
+We delved deeper into this phenomenon and identified the “shortcut learning problem” [12] inherent in lip sync task.
+The shortcut learning problem in lip sync.
+Lip-sync methods are typically based on a video-to-video inpainting framework, the model receives masked frames and audio as inputs.
+Unexpectedly, the audio-conditioned LDMs tends to predict lip movements based on visual information around the lips, such as facial muscles, eyes, and cheeks, while ignoring the audio information.
+We conducted an experiment to validate the existence of the shortcut learning problem and the effectiveness of SyncNet supervision [29] in mitigating this issue.
+We trained the audio-conditioned LDMs with masks of different sizes, both with and without SyncNet supervision.
+We used the sync confidence score [9] to evaluate the synchronization accuracy between audio and lip movements.
+Figure 2: The shortcut learning problem in the lip-sync task. Higher sync confidence score means better lip-sync accuracy.
+As shown in Fig.
+2, without SyncNet supervision, as the mask size increases, the visual information available in the masked frame for inferring lip movements decreases, forcing the model to rely more on audio information, thereby improving lip-sync accuracy, and vice versa.
+In contrast, with SyncNet supervision, the model remains focused on learning audio-visual correlations across different mask sizes.
+Why is SyncNet supervision not necessary for audio-driven portrait animation methods?
+These methods [41, 46, 7, 20] are based on image-to-video framework.
+Since they do not involve input masked frames, they do not suffer from the shortcut learning problem.
+The previous work Diff2Lip [28] has explored how to add SyncNet supervision to pixel-space diffusion models.
+However, how to effectively apply SyncNet to latent diffusion models [33] remains unclear.
+Specifically, we explored two methods to incorporate SyncNet supervision into latent diffusion models: (a) Decoded pixel space supervision and (b) Latent space supervision.
+Furthermore, We found that SyncNet struggles to converge in both latent space and high-resolution pixel space.
+Since the convergence of SyncNet significantly impacts its supervision effectiveness, and ultimately affects the lip-sync accuracy of the supervised model, the convergence issue of SyncNet is highly valuable for research.
+Therefore, we conducted the first comprehensive empirical studies in the aspects of model architecture design, training hyperparameters, and data preprocessing techniques, introducing the StableSyncNet with an architecture designed for stable convergence.
+Our StableSyncNet achieved the unprecedented 94% accuracy on HDTF [51].
+Additionally, we observed that high-frequency details in the generated talking videos, such as teeth, lips, and facial hair, exhibit flickering artifacts.
+We propose TREPA, a novel method designed to enhance temporal consistency and reduce such artifacts.
+In summary, we made the following contributions: (1) We proposed LatentSync, the first lip-sync method that utilizes audio-conditioned LDMs to achieve end-to-end lifelike lip sync on high-resolution videos, incorporating TREPA to enhance the temporal consistency in the generated videos.
+(2) We identified the shortcut learning problem in lip-sync task and explored different methods to incorporate SyncNet supervision into audio-condtioned LDMs.
+(3) We conducted the first comprehensive empirical studies to identify key factors affecting SyncNet convergence, introducing the StableSyncNet for stable convergence.
 Figure 3: The overview of our LatentSync framework.
@@
 The reference and masked frames are channel-wise concatenated with noised latents as the input of U-Net.
-In the training process, we use one-step method to get estimated clean latents from predicted noises, which are then decoded to obtain the estimated clean frames.
+In the training process, we use a one-step method to get estimated clean latents from predicted noises, which are then decoded to obtain the estimated clean frames.
 The TREPA, LPIPS [49] and SyncNet loss [29] are added in the pixel space.
+2 Related Work
+2.1 Diffusion-based Lip Sync
+Diff2Lip [28] and DrivenVideoEditing [3] are both end-to-end lip sync methods based on pixel-space audio-conditioned diffusion models.
+MyTalk [48] uses diffusion models in the first stage to complete audio-to-motion conversion and uses a VAE [23] in the second stage for motion-to-image generation.
+StyleSync [53] uses transformers in the first stage to convert audio to motion and employs diffusion models in the second stage for motion-to-image generation.
+DiffDub [25] uses diffusion autoencoders [30] to convert the masked images into semantic latent codes in the first stage, and uses diffusion models to generate an image conditioned on the semantic latent codes and audio in the second stage.
+2.2 Non-diffusion-based Lip Sync
+Wav2Lip [29] is the most classic lip sync method that introduced using a pretrained SyncNet [9] to supervise the training of the lip-sync generator.
+[16] trains a VQ-VAE [11] to encode faces and head poses, and then trains the lip sync generator in the quantized space to generate high-resolution images.
+StyleSync [14] follows the overall framework of Wav2Lip, with its main innovation being the use of StyleGAN2 [22] as the generator backbone.
+VideoReTalking [8] divides lip sync into three components: semantic-guided reenactment network, lip sync network, and identity-aware refinement and enhancement.
+DINet [52] deforms the feature maps to generate mouth shapes conditioned on the driving audio.
+MuseTalk [50] uses the architecture of Stable Diffusion [33] for inpainting, but it does not perform diffusion process and uses a discriminator for adversarial learning [13], making it more like a GAN-based framework.
+2.3 Audio-driven Portrait Animation
+Many people may confuse lip sync with audio-driven portrait animation.
+These two tasks have some similarities but are actually completely different tasks.
+Lip sync is based on video-to-video editing framework, which needs to keep areas other than the mouth the same as in the input video.
+Audio-driven portrait animation is based on image-to-video animation framework, which can change the head movement and even facial expressions.
+Although there are already some audio-driven portrait animation methods based on audio-conditioned LDMs, such as [41, 7, 46, 20], they cannot be directly applied to the lip sync task due to the shortcut learning problem.
+3 Method
 3.1 LatentSync Framework
 The overview of the LatentSync framework is shown in Fig.
-2.
-Our backbone network generates the video frame by frame without temporal modeling between frames.
-The temporal consistency during the generation process relies entirely on the audio window.
-This frame-by-frame generation approach is beneficial for occlusion strategies in dubbing scenarios, as we can easily skip frames where the mouth is occluded and lip sync is not needed.
-Our framework is based on image-to-image inpainting [43], so it requires a masked image as input.
-Additionally, to integrate the visual features of the face from the original video, we need to input reference images.
-The reference images are randomly selected during training to avoid learning shortcuts [16], and they are taken from the current frames during inference.
-We concatenate these inputs along the channel dimension, so the total input consists of 13 channels (4 channels for noise latent, 1 channel for the mask, 4 channels for the masked image, and 4 channels for the reference image).
-At the beginning of training, we initialized our model with the parameters of SD 1.5 [41], except for the first conv_in layer with 13 channels and cross-attention layers of dimension 384, which are randomly initialized.
-Theoretically, our framework is compatible with more advanced latent diffusion models such as SDXL [35], SD3 [13], and FLUX [28], but we only conducted experiments on SD 1.5 due to hardware limitations.
+3.
+The framework is based on video-to-video inpainting with the temporal modeling by temporal layer [15].
+To incorporate the visual features of the face from the input video, reference frames are introduced as additional inputs.
+During training, these frames are randomly selected, while during inference, they are taken from the current frames.
+We concatenate different inputs along the channel dimension, making the total input of U-Net to be 13 channels (4 channels for noise latent, 1 channel for the mask, 4 channels for the masked frame, and 4 channels for the reference frame).
+At the beginning of training, the model is initialized with the parameters of SD 1.5 [33], except for the first conv_in layer with 13 channels and cross-attention layers of dimension 384, which are randomly initialized.
 Audio layers.
@@
 To integrate the audio embeddings into the U-Net [34], we used the native cross-attention layer.
+Affine transformation and fixed mask.
+During the data preprocessing stage, affine transformation was employed to perform face frontalization.
+This approach [14] helps the model to effectively learn facial features particularly in challenging scenarios such as side-profile views.
+We applied a mask that covers the entire face to minimize the model’s tendency to learn visual-visual shortcuts.
+The position and shape of the mask are fixed.
+We do not use the detected landmarks [26, 5] to draw the mask, as moving landmarks also provide cues about lip movements.
+The affine transformation and fixed mask are illustrated in Fig.
+4.
+Figure 4: The illustration of affine transformation and fixed mask.
 SyncNet supervision.
-Our model predicts in the noise space, while SyncNet [9, 36] requires an input in the image space.
+Latent diffusion models predict in the noise space, while SyncNet [9, 29] requires an input in the image space.
 To address this issue, we use the predicted noise $\epsilon_{\theta}(z_{t})$ to obtain the estimated $\hat{z}_{0}$ in one step, which can be formulated as:
 $\hat{z}_{0}=\left(z_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(z_{t})\right)/\sqrt{\bar{\alpha}_{t}}$ (1)
-Another problem is that our model makes predictions in the latent space.
-We explored two methods to add SyncNet supervision to latent diffusion models: (a) Decoded pixel space supervision, which trains SyncNet in the same way as [36].
+Another problem is that latent diffusion models make predictions in the latent space.
+We explored two methods to incorporate SyncNet supervision into latent diffusion models: (a) Decoded pixel space supervision, which trains SyncNet in the same way as Wav2Lip [29].
 (b) Latent space supervision, which requires training a SyncNet in the latent space.
@@
 The illustration is shown in Fig.
-3.
-We found that training SyncNet in the latent space converges worse compared to training it in the pixel space.
-We speculate that this is due to information loss in the lip area during the VAE encoding process.
-The poorer convergence of the latent space SyncNet also results in the worse audio-lip sync accuracy for the supervised diffusion models, according to the experimental results in Sec.
-4.3.
-Therefore, we use decoded pixel space supervision in the LatentSync framework.
+5.
+Our empirical analysis in Sec.
+5.3 reveals that training SyncNet in the latent space exhibits inferior convergence compared to training in the pixel space.
+This degradation may arise from information loss in the lip region during the VAE encoding process.
+The poorer convergence of SyncNet in the latent space adversely impacts the lip-sync accuracy of the supervised diffusion models, according to the experimental results in Sec.
+5.3.
+Therefore, we finally choose the decoded pixel space supervision in the LatentSync framework.
 Figure 5: Two methods to add SyncNet supervision to latent diffusion models.
-Two-stage training.
-In the first stage, we do not decode to the pixel space and do not add SyncNet loss.
-The purpose of the first stage is to allow the U-Net to learn visual features using a large batch size.
-Therefore, the training objective of the first stage has only a simple loss [23]:
+3.2 Two-Stage Training Strategy
+The decoded pixel space supervision has a problem that the activations in VAE decoding need to be stored for backpropagation, which significantly increases GPU memory consumption.
+To mitigate this issue, we designed a two-stage training strategy: in the first stage, the model learns to extract features from reference frames and develops inpainting capabilities, which we refer to as learning visual features.
+In the second stage, it learns audio-visual correlations under SyncNet supervision.
+This approach eliminates the need for the VAE decoding process in the first stage, enabling model to learn visual features with a larger batch size.
+We observed that audio-conditioned LDMs typically spend more time learning visual features and less time learning audio-visual correlations.
+Therefore, this two-stage training strategy allows the model to efficiently learn both visual features and audio-visual correlations.
+We provide the formal definitions of training objectives in the following paragraphs.
+In the first stage of training, we do not add the temporal layer [15, 4] and train all parameters of the U-Net. The training objective has only a simple loss [19]:
 $\mathcal{L}_{\text{simple}}=\mathbb{E}_{x,A,\epsilon\sim\mathcal{N}(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(A))\right\|_{2}^{2}\right]$ (2)
 where $A$ is the input audio, $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(A))$ is the predicted noise, and $\tau_{\theta}$ is the audio feature extractor.
-In the second stage of training, we used the decoded pixel space supervision method to add SyncNet loss.
-Compared to [36], we increased the input frame length of SyncNet to 16.
-For detailed changes to the SyncNet structure, please refer to Sec.
-3.3.
-By using the 16 decoded video frames $\mathcal{D}(\hat{z}_{0})_{f:f+16}$ and the corresponding audio sequence $a_{f:f+16}$, the SyncNet loss can be formulated as:
+In the second stage of training, we only train the temporal layer and audio layer while freezing the other parameters of the U-Net.
+Suppose we have 16 decoded video frames $\mathcal{D}(\hat{z}_{0})_{f:f+16}$ and the corresponding audio sequence $a_{f:f+16}$, the SyncNet loss can be formulated as:
 $\mathcal{L}_{\text{sync}}=\mathbb{E}_{x,a,\epsilon,t}\left[\text{SyncNet}(\mathcal{D}(\hat{z}_{0})_{f:f+16},a_{f:f+16})\right]$ (3)
@@
 In addition, to improve temporal consistency, we also employed the proposed TREPA, see Eq.
-7.
+6.
 For more details of the TREPA, please refer to Sec.
-3.2.
+3.3.
 The total loss function for the second stage of training is:
 $\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{\text{simple}}+\lambda_{2}\mathcal{L}_{\text{sync}}+\lambda_{3}\mathcal{L}_{\text{lpips}}+\lambda_{4}\mathcal{L}_{\text{trepa}}$ (5)
-We identified the optimal coefficients through empirical analysis as follows: $(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})=(1,0.05,0.1,10)$
-Mixed noise model.
-To ensure that the model correctly learns temporal information, the input noises also need to have temporal consistency; otherwise the model will fail to learn properly.
-The simplest method is applying the same noise to all frames [34].
-We further found that diffusion models trained with the mixed noise model [14] exhibit better temporal consistency.
-The method generates two distinct noise vectors, $\epsilon_{\text{shared}}$ and $\epsilon_{\text{ind}}$.
-$\epsilon_{\text{shared}}$ represents a common noise vector applied across all 16 video frames, whereas $\epsilon_{\text{ind}}$ corresponds to the frame-specific noise.
-The final noise is obtained through a linear combination of these two vectors.
-$\epsilon^{f}=\epsilon_{\text{shared}}+\epsilon_{\text{ind}}^{f}$ (6)
-We used the mixed noise model in both two training stages, and we applied the same initial latent noise to all frames in the video during inference stage.
-Affine transformation.
-During the data preprocessing stage, affine transformation was employed to achieve face frontalization.
-This approach [18] helps the model to effectively learn facial features particularly in challenging scenarios such as side-profile views.
-We applied a mask that covers the entire face to prevent information leakage, ensuring that the model does not infer lip movements based on facial movements.
-The position and shape of the mask are fixed.
-We do not use detected landmarks [30, 5] to draw the mask as they will leak information about facial movements.
-The affine transformation and fixed mask are illustrated in Fig.
-4.
-Figure 4: The illustration of affine transformation and fixed mask.
 3.3 Temporal Representation Alignment
 TREPA aligns the temporal representations of the generated image sequences with those of ground truth image sequences.
-The insight behind this method is that merely employing distance loss between individual images only improves the content quality of single generated images but does not enhance the temporal consistency of the generated image sequence.
+The insight behind this method is that merely employing distance loss between individual images improves the content quality of single generated images but does not enhance the temporal consistency of the generated image sequence.
 In contrast, temporal representations capture temporal correlation within image sequences, enabling the model to focus on improving the overall temporal consistency.
@@
 We also fix the representation by $\ell_{2}$ normalization before calculating the MSE.
-Observations.
-The temporal layer [19, 4] adds temporal self-attention layers to the U-Net to enhance temporal consistency.
-We initially attempted to apply the temporal layer to LatentSync, but found that it significantly compromised lip-sync accuracy.
-In contrast, our TREPA not only avoids harming lip-sync accuracy but even improves it to some extent.
-We hypothesize this is because the temporal layer adds additional parameters to the U-Net, causing some of the gradients in backpropagation to be distributed to the temporal layer parameters, weakening the learning of the audio cross-attention layer parameters.
-On the other hand, TREPA does not add extra parameters; to improve temporal consistency, the model must rely more effectively on the information in the audio window (since we know that lip-sync models use the audio window to capture temporal information).
-As a result, the audio cross-attention layers are further trained and strengthened.
-Refer to the quantitative results in Sec.
-4.3.
-Another interesting finding is that TREPA can also be applied to improve the temporal consistency and lip-sync accuracy of Wav2Lip [36], according to the experimental results in Sec.
-4.3.
-Theoretically, any video generation model that generates multiple consecutive frames at once can utilize TREPA to enhance the temporal consistency.
-3.3 Empirical Studies on SyncNet Convergence
-Many researchers have reported the convergence problem of SyncNet on GitHub issues [32, 56, 40, 33], a typical characteristic of this issue is that the training loss gets stuck at 0.69 and fails to decrease further.
-However, no academic work has provided a detailed analysis of this problem to date.
-In this section, we analyze the SyncNet convergence problem and identify several critical factors affecting the convergence of SyncNet through various ablation studies.
-Importantly, we preserved SyncNet’s original training framework and employed the same contrastive loss function as the original paper [9].
-Therefore, our experience can be applied to many lip sync [36, 18, 34] and audio-driven portrait animation methods [31] that utilize SyncNet.
+4 Empirical Studies on SyncNet Convergence
+Our experiments reveal that SyncNet is difficult to converge in both latent space and high-resolution pixel space, a typical characteristic of this issue is that the training loss gets stuck at 0.69 and fails to decrease further.
+In this section, we analyze the SyncNet convergence problem and identify several critical factors affecting the convergence through various ablation studies.
+Importantly, we preserved SyncNet’s original training framework and employed the same contrastive loss function as Wav2Lip [29].
+Therefore, our experience can be applied to many lip-sync [29, 14, 28, 8, 52, 50] and audio-driven portrait animation methods [27, 47] that utilize SyncNet.
 Why is the SyncNet training loss stuck at 0.69?
-According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage.
+According to the classic training framework of SyncNet [29, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage.
 The output of SyncNet is a probability distribution of whether the sample is positive or negative.
@@
 Let $p(x)$ be the true probability distribution and $q(x)$ the predicted probability distribution.
-When $q(x=1)\approx q(x=0)\approx 0.5$ and batch size $N$ is sufficiently large, the training loss of SyncNet is (proof in Appendix B):
+When $q(x=1)\approx q(x=0)\approx 0.5$ and batch size $N$ is sufficiently large, the training loss of SyncNet is (proof in Appendix A):
 $\mathcal{L}_{\text{syncnet}}$ $=-\frac{1}{N}\sum_{i=1}^{N}\sum_{x_{i}\in\{0,1\}}p(x_{i})\,\text{log}\,{q(x_{i})}\approx 0.693$ (7)
@@
 As shown in Fig.
-5, a larger batch size (e.g., 1024) not only enables the model to converge faster and more stably but also results in a lower validation loss at the end of training.
+6, a larger batch size (e.g., 1024) not only enables the model to converge faster and more stably but also results in a lower validation loss at the end of training.
 In contrast, smaller batch sizes (e.g., 128) may fail to converge, with the loss remaining stuck at 0.69.
@@
 VoxCeleb2 results, the more transparent curves represent the training set loss, while the darker curves represent the validation set loss; the same applies to the following figures.
-(VoxCeleb2, Dim 2048, SD residual blocks, 16 frames.)
+(VoxCeleb2, Dim 2048, StbleSyncNet arch, 16 frames.)
 Architecture.
-We attempted to replace the SyncNet’s visual and audio encoders with the U-Net encoder from SD 1.5 [41], retaining the structure of residual blocks [21] and self-attention blocks [51] in the U-Net encoder blocks.
+We redesign the SyncNet’s visual and audio encoders with the U-Net encoder from Stable Diffusion 1.5 [33], retaining the structure of residual blocks [17] and self-attention blocks [43] in the U-Net encoder blocks.
 We only adjusted the downsampling factors based on the size of the input visual images and mel-spectrograms, and we removed the cross-attention blocks, as SyncNet does not require additional conditions.
+We refer to the SyncNet with this modified architecture as StableSyncNet.
 As shown in Fig.
-6, SyncNet using the SD U-Net encoder architecture maintained both training loss and validation loss lower than those of Wav2Lip’s SyncNet [36] throughout the training process, though the difference was not substantial.
+7, StableSyncNet maintained both training loss and validation loss lower than those of Wav2Lip’s SyncNet [29] throughout the training process.
 Figure 7: SyncNet training curves of different architectures.
@@
 (VoxCeleb2, Dim 2048, Batch size 512, 5 frames.)
-Self attention.
-We attempted to remove the self-attention [51] blocks in the U-Net encoder, keeping only the residual blocks [21].
-As shown in Fig.
-7, removing the self-attention blocks results in faster convergence during the initial training phase, while achieving lower validation loss towards the end of the training.
-Additionally, this approach helps to reduce the model size and accelerates both inference and training speeds.
-We hypothesize that the reason may be that attention layers do not possess the inherent inductive biases of CNNs, such as translation equivariance and locality.
-Therefore, they do not generalize well without sufficient data.
-Similar findings and explanations can be found in [11].
-Figure 7: SyncNet training curves with and without self-attention layers. (VoxCeleb2, Dim 2048, Batch size 512, 16 frames.)
+Embedding dimension.
+As illustrated in Fig.
+8, embeddings with smaller dimensions (e.g., 512) result in representations that fail to capture sufficient semantic information, while larger dimensions (e.g., 4096 or 6144) lead to sparse representations, thereby impeding model convergence.
+Se identified that an optimal embedding dimension of 2048 is suitable for images with an input resolution of 256 $\times$ 256 pixels.
+Figure 8: SyncNet training curves of different embedding dimensions. (VoxCeleb2, Batch size 512, StableSyncNet arch, 16 frames.)
 Number of frames.
@@
 As shown in Fig.
-8, selecting a larger number of frames (e.g., 16) can help the model converge.
+9, selecting a larger number of frames (e.g., 16) can help the model converge.
 However, an excessively large number of frames (e.g., 25) will cause the model to get stuck around 0.69 in the early stage of training, and the validation loss at the end of training does not show a significant advantage compared to using 16 frames.
-Figure 8: SyncNet training curves of different numbers of input frames. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048.)
+Figure 9: SyncNet training curves of different numbers of input frames. (VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 2048.)
 Data preprocessing.
-In-the-wild videos naturally contain audio-visual offsets, so it is necessary to adjust this offset to zero before inputting them into the SyncNet network.
-We used the official open-source version of SyncNet [9] to adjust the offset and remove videos with ${\text{Sync}}_{\text{conf}}$ below 3.
-Specifically, we tested adjusting the offset before and after applying affine transformation.
+In-the-wild videos naturally contain audio-visual offsets, it is necessary to adjust this offset to zero before inputting them into the SyncNet network.
+We used the official open-source version of pretrained SyncNet [9] to adjust the offset and remove videos with ${\text{Sync}}_{\text{conf}}$ below 3.
+Specifically, we evaluated adjusting the offset before and after applying affine transformation.
 As shown in Fig.
-9, without offset adjustment, the model’s convergence is significantly impaired.
+10, without offset adjustment, the model’s convergence is significantly impaired.
 Performing affine transformation before offset adjustment yields better results.
-We hypothesize that this may be due to affine transformation reducing data with side profiles or unusual angles, allowing the official SyncNet [9] to predict the offset more accurately.
-Figure 9: SyncNet training curves of different data preprocessing methods. (VoxCeleb2, Batch size 512, SD residual blocks, Dim 2048, 16 frames.)
+This may be due to that affine transformation reducing data with side profiles or unusual angles, allowing the pretrained SyncNet [9] to predict the offset more accurately.
+Figure 10: SyncNet training curves of different data preprocessing methods. (VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 2048, 16 frames.)
 Discussions.
-Based on the ablation studies above, we found that batch size, number of input frames, and data preprocessing method significantly impact SyncNet convergence.
-The model architecture can also affect convergence, though less noticeably.
-We identified the optimal settings for a SyncNet with the 256 $\times$ 256 input size: batch size of 1024, 16 frames, SD residual blocks for both visual and audio encoders, embedding dimension of 2048, and adjusting the audio-visual offset after affine transformation.
-We train a SyncNet on VoxCeleb2 [10] and test it on HDTF [62], which is an out-of-distribution experimental setup.
+Based on the ablation studies above, we found that batch size, number of input frames, and data preprocessing method are the primary factors affecting SyncNet convergence.
+We identified the optimal settings for a StableSyncNet with the 256 $\times$ 256 input size: batch size of 1024, 16 frames, SD U-Net encoder adapted for both visual and audio encoders, embedding dimension of 2048, and adjusting the audio-visual offset after affine transformation.
+We train a StableSyncNet on VoxCeleb2 [10] and test it on HDTF [51], which is an out-of-distribution experimental setup.
 The validation loss on VoxCeleb2 reaches around 0.18 and the accuracy on HDTF achieves 94%, which significantly surpasses the previous SOTA result 91% [29, 16].
-For more ablation studies, please refer to Appendix C.
 5 Experiments
 Method HDTF VoxCeleb2
-FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
-Wav2Lip [36] 12.5 0.70 8.2 304.35 10.8 0.71 7.0 257.85
-VideoReTalking [8] 9.5 0.75 7.5 270.56 7.5 0.77 6.4 215.67
-DINet [63] 8.02 0.79 7.7 208.93 6.3 0.80 6.7 110.75
-MuseTalk [61] 9.35 0.74 6.8 246.75 7.1 0.80 5.9 203.43
-LatentSync (Ours) 7.03 0.79 8.9 192.74 5.6 0.81 7.3 124.38
+FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow LMD $\downarrow$ FVD $\downarrow$ FID $\downarrow$ SSIM $\uparrow$ ${\text{Sync}}_{\text{conf}}$uparrow LMD $\downarrow$ FVD $\downarrow$
+Wav2Lip [29] 12.5 0.70 8.2 0.34 304.35 10.8 0.71 7.0 0.53 257.85
+VideoReTalking [8] 9.5 0.75 7.5 0.49 270.56 7.5 0.77 6.4 0.60 215.67
+Diff2Lip [28] 10.3 0.72 7.9 0.36 260.45 9.8 0.73 6.9 0.54 210.45
+MuseTalk [50] 9.35 0.74 6.8 0.56 246.75 7.1 0.80 5.9 0.64 203.43
+LatentSync (Ours) 7.22 0.79 8.9 0.30 162.74 5.7 0.81 7.3 0.51 123.27
 Table 1: Quantitative comparisons on HDTF and VoxCeleb2.
+Figure 11: Qualitative comparisons with SOTA lip-sync methods.
+We run two cases in the cross generation setting [28].
+The first row demonstrates the original input video, and the second row is the video from which we extracted the audio as input, the video can be regarded as the target lip movements.
+Rows 3 $\sim$ 7 display the lip-synced videos.
+(All the photorealistic portrait images in this paper are from contracted models.)
 5.1 Experimental Settings
@@
 (2) Lip-sync accuracy.
-We use the confidence score of SyncNet (${\text{Sync}}_{\text{conf}}$) [9], which we found to align closely with visual assessments.
+We use the confidence score of SyncNet (${\text{Sync}}_{\text{conf}}$) [9] and the landmark distances around the mouth (LMD) [14].
+We found that ${\text{Sync}}_{\text{conf}}$ aligns closely with visual assessments.
 (3) Temporal consistency.
@@
 VideoReTalking [8] divides the lip-sync process into three steps to improve the results.
-DINet [63] proposed performing deformations on feature maps to achieve high-quality lip sync.
+Diff2Lip [28] utilized pixel-space diffusion models to achieve generalized lip sync.
 MuseTalk [50] utilizes the architecture of Stable Diffusion for inpainting, but it is not based on the diffusion model framework and appears more like a GAN-based approach.
@@
 1, the lip-sync accuracy of our method significantly surpasses that of other methods.
-We believe this may be attributed to our 94% accuracy SyncNet, as well as the audio cross-attention layers, which may better capture the relationship between audio and lip movements.
-In terms of visual quality, our method outperforms others, likely due to the capabilities of the Stable Diffusion model.
-Owing to the incorporation of TREPA, our FVD score is also superior to other methods.
+This is attributed to our 94% accuracy StbaleSyncNet, as well as the audio cross-attention layers in U-Net, which better captures the relationship between audio and lip movements.
+In terms of visual quality, our method outperforms others, likely due to the powerful capabilities of the Stable Diffusion model.
+Owing to temporal layer [15] and the incorporation of TREPA, our FVD score is also superior to other methods.
 Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
-LatentSync w/o SyncNet 1.6 260.37
-LatentSync + latent space SyncNet 7.9 240.45
-LatentSync + pixel space SyncNet 8.9 192.74
+LatentSync w/o SyncNet 4.6 220.37
+LatentSync + latent space SyncNet 7.9 180.45
+LatentSync + pixel space SyncNet 8.9 162.74
 Table 2: The ablation studies of different SyncNets. HDTF results.
@@
 According to Fig.
-10, Wav2Lip has excellent lip-sync accuracy, but the generated videos are very blurry.
+11, Wav2Lip has excellent lip-sync accuracy, but the generated videos are very blurry.
 VideoReTalking shows strange artifacts.
-DINet’s generated box is obvious, and its mouth movements are stiff.
+Diff2Lip is limited by pixel-space diffusion models and can only generate low-resolution videos, resulting in noticeable blurriness.
 MuseTalk does not preserve facial features well, the man’s beard becomes sparse.
@@
 Furthermore, it hardly shows generated box due to the smooth shape of fixed mask.
-Figure 10: Qualitative comparisons with SOTA lip-sync methods.
-We run two cases in the cross generation setting [34].
-The first row demonstrates the original input video, and the second row is the video from which we extracted the audio as input, the video can be regarded as the target lip movements.
-Rows 3 $\sim$ 7 display the lip-synced videos.
-(All the photorealistic portrait images in this paper are from licensed models.)
 Figure 12: SyncNet training curves of different visual input spaces.
 Here we use the open-sourced pretrained VAE from Stability AI [1] for encoding.
-(VoxCeleb2, Batch size 512, SD residual blocks, Dim 1024, 16 frames.)
+(VoxCeleb2, Batch size 512, StableSyncNet arch, Dim 1024, 16 frames.)
 5.3 Ablation studies
-The supervision effect of SyncNet.
+The effectiveness of SyncNet supervision.
 As shown in Tab.
-2, when the SyncNet supervision is not added, the trained diffusion models have no lipsync ability at all.
+2, when the SyncNet supervision is not added, The lip-sync performance of the trained diffusion models is significantly poor.
 In fact, we found that other end-to-end lipsync methods [29, 28] exhibit similar phenomena.
 As for supervision in two different spaces, diffusion models under pixel space SyncNet supervision perform better in terms of lip-sync accuracy and temporal consistency.
-This is likely due to the poor convergence of latent space SyncNet supervision (as shown in Fig.
-11), which is quite reasonable since the input to the latent space SyncNet is the compressed latents obtained from VAE encoding, and some lip information may already be lost.
+This is likely due to the poor convergence of latent space SyncNet (as shown in Fig.
+12), which is reasonable since the input to the latent space SyncNet is the compressed latents obtained from VAE encoding, and some lip information may already be lost.
 In addition, we found that improvements in lip-sync accuracy were accompanied by an increase in temporal consistency.
-This is because the temporal consistency of our model is provided by the audio window; as the model becomes better at utilizing audio information, temporal consistency also improves accordingly.
-Different methods for improving temporal consistency.
+This is because the audio window inherently encapsulates rich temporal information.
+Enhanced utilization of audio information by the model also leads to the improved temporal consistency.
+The effectiveness of TREPA.
 According to Tab.
-3, the temporal layer significantly improves temporal consistency but also noticeably harms lip-sync accuracy.
-Our TREPA achieves balanced improvements in both lip-sync accuracy and temporal consistency without evident drawbacks.
-Considering the importance of lip-sync accuracy, TREPA is obviously the better approach.
-Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
-LatentSync 8.2 234.35
-LatentSync + Temporal layer [19] 4.7 174.56
-LatentSync + TREPA 8.9 192.74
-Table 3: The ablation studies of different methods for improving temporal consistency. HDTF results.
-In addition, TREPA can also improve Wav2Lip according to Tab. 4, which suggests that our method has general applicability in lip-sync tasks.
-Method ${\text{Sync}}_{\text{conf}}$uparrow FVD $\downarrow$
-Wav2Lip 8.2 304.35
-Wav2Lip + TREPA 8.6 264.56
-Table 4: The ablation studies of Wav2Lip with TREPA. HDTF results.
+3, both the temporal consistency and visual quality improve after incorporating TREPA.
+This improvement may be attributed to that the robust representations extracted by VideoMAE-v2 [44] inherently encapsulate both visual and temporal information.
+Method FID $\downarrow$ SSIM $\uparrow$ FVD $\downarrow$
+LatentSync 7.71 0.77 176.35
+LatentSync + TREPA 7.22 0.79 162.74
+Table 3: The ablation studies of TREPA. HDTF results.
 6 Conclusion
-We introduced a new lip sync framework named LatentSync that addresses the problems of traditional diffusion-based lip sync methods: (1) The low sampling speed and inability to generate high-resolution videos of pixel space diffusion.
+We introduced LatentSync, the first lip-sync method based on audio-conditioned LDMs, which addresses the problems of traditional diffusion-based lip-sync methods: (1) The low sampling speed and inability to generate high-resolution videos of pixel space diffusion.
 (2) The information loss in two-stage generation methods.
-In addition, we proposed TREPA to improve temporal consistency while preserving the lip-sync accuracy, which mitigates the problem of video flickering in diffusion-based lip sync methods.
-Furthermore, we analyzed the notorious SyncNet convergence issue and conducted comprehensive ablation studies to identify several key influencing factors.
+We identified the shortcut learning problem in lip-sync task and explored different approaches to incorporate SyncNet supervision to solve this problem.
+We conducted comprehensive ablation studies to find out several key factors influencing SyncNet converge.
+In addition, we proposed TREPA to further improve the temporal consistency of our method.
 Appendix
-Appendix A The Differences between Lip Sync and Audio-driven Portrait Animation
-Many people may confuse lip sync with audio-driven portrait animation. These two tasks have some similarities but are actually completely different tasks. Their main difference lies in the input:
-  • The inputs of Lip Sync: Video + Audio
-  • The inputs of Audio-driven Portrait Animation: Image + Audio
-This leads to differences in their overall frameworks.
-Lip sync is more like a video-to-video editing framework, which needs to keep areas other than the mouth the same as in the input video.
-Audio-driven portrait animation is more like an image-to-video animation framework, which can change the head’s movement and even facial expressions.
-The differences in the overall frameworks result in substantial differences in their training and inference methods.
-We list some methods for the two different tasks to help readers better distinguish between them:
-  • Lip Sync: Wav2Lip[36], Diff2Lip [34], StyleSync [18], MuseTalk [61], MyTalk [57]
-  • Audio-driven Portrait Animation: EMO [49], Hallo [54], EchoMimic [7], VASA-1 [55], DreamTalk [31], SadTalker [60]
-Although there are already some audio-driven portrait animation methods based on Stable Diffusion [41], such as EMO [49] and Hallo [54], it is hard to adapt them for the lip sync task.
-For example, the temporal layer is very effective in these methods, but we found that it is not suitable for the lip sync task.
-To the best of our knowledge, the methods closest to ours are Diff2Lip [34] and MyTalk [57], which are pixel space diffusion and two-stage generation methods.
-Appendix B SyncNet Training Loss Proof
-According to the classic framework of SyncNet [36, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage.
+Appendix A SyncNet Training Loss Proof
+According to the classic training framework of SyncNet [29, 9], we randomly provide SyncNet with positive and negative samples with a 50% probability during the training stage.
 The output of SyncNet is a probability distribution of whether the sample is positive or negative.
@@
 As shown in Fig.
-12, in the early stages of training, the probabilities predicted by SyncNet for $q(x_{i}=1)$ are dispersed, but later in the training, the probabilities for $q(x_{i}=1)$ are almost all distributed around 0.5 (the scatter charts for $q(x_{i}=0)$ are similar).
+13, in the early stages of training, the probabilities predicted by SyncNet for $q(x_{i}=1)$ are dispersed, but later in the training, the probabilities for $q(x_{i}=1)$ are almost all distributed around 0.5 (the scatter charts for $q(x_{i}=0)$ are similar).
 Now we have:
@@
 $=0.693$ (15)
-Appendix C Additional SyncNet Ablation Studies
-C.1 Embedding dimension
-The dimensions of the audio and visual embeddings in SyncNet should be carefully calibrated to a moderate size.
-As illustrated in Fig.
-13, embeddings with smaller dimensions (e.g., 512) result in representations that fail to capture sufficient semantic information, while larger dimensions (e.g., 4096 or 6144) lead to sparse representations, thereby impeding model convergence.
-Through empirical analysis, we identified that an optimal embedding dimension of 2048 is suitable for images with an input resolution of 256 $\times$ 256 pixels.
-For images of varying resolutions, the embedding dimensions should be appropriately scaled to ensure optimal performance.
-Figure 13: SyncNet training curves of different embedding dimensions. (VoxCeleb2, Batch size 512, SD residual blocks, 16 frames.)